OpenStudy (anonymous):

Prove that there exists no natural numbers , n and n such that m^2 =n^2 +2002

Miracrown (miracrown):

I would compare this to the pythagorean theorem

Miracrown (miracrown):

so 2002 would need to be a perfect square and it is not

Miracrown (miracrown):

for example we can have 5^2 = 3^2 + 4^@

OpenStudy (anonymous):

\[3^2=2^2+5\]

Miracrown (miracrown):

then in order for m and n to be natural numbers 2002 would have to be a perfect square and it clearly isn't

OpenStudy (anonymous):

and 5 is not a perfect square

Miracrown (miracrown):

5^2 is a perfect square

OpenStudy (anonymous):

My teacher gave a hint to make cases like , but after that i don't know what to do , 

OpenStudy (anonymous):

@Miracrown what i said was \[3^2=2^2+5\] and \(5\) is not a prefect square

OpenStudy (anonymous):

think this has to do with remainders of perfect squares

Miracrown (miracrown):

they want us to compare what happens when m and n are odd and even

Miracrown (miracrown):

so the idea is if m is odd then m^2 has to be odd also

OpenStudy (anonymous):

(m+n)(m-n)=2002 It can be represented like this

Miracrown (miracrown):

ok do you need to go through the algebra for that step?

OpenStudy (raden):

m^2 - n^2 = 2002 (m+n)(m-n) = factors of 2002 m+n = 2002 m-n = 1 m+n = 1001 m-n = 2 m+n = 286 m-n = 7 m+n = 182 m-n = 11 m+n = 154 m-n = 13 m+n = 143 m-n = 14 m+n = 91 m-n = 22 m+n = 77 m-n = 26 all case above solve by elimination or subtituion, obvious m is not a natural number

OpenStudy (anonymous):

looks good to me!

OpenStudy (anonymous):

yes neat thanks

OpenStudy (zarkon):

using odd and even you can write a more elegant proof in just a few lines

OpenStudy (anonymous):

How

OpenStudy (anonymous):

(m+n) *(m-n)=2002 since 2002 is even number if either m or n is odd then m+n and m-n are both odds and hence their product must be odd but 2002 is even if both m and n are even then 2|m+n and 2|m-n as well and hence (m+n)*(m-n) must be divisible by 4 but 2002 is not divisible by 4

OpenStudy (zarkon):

for even assume n,m are even then m=2k and n=2l then \[(2k)^2=(2l)^2+2002\] \[4k^2=4l^2+2002\] \[2k^2-2l^2=1001\] the lhs is even the rhs is odd a contadiction

OpenStudy (anonymous):

similarly if m and n are both odd then 2|m+n and 2|m-n as well and hence (m+n)*(m-n) must be divisible by 4 but 2002 is not divisible by 4 so the result follows...